OpenStudy (anonymous):

The side length of a rectangular box with a square base is increasing at the rate of 2 ft/sec, while the height is decreasing at the rate of 2 ft/sec. At what rate is the volume of the box changing when the side length is 10 ft and the height is 12 ft? (5 points) 680 ft3/sec 40 ft3/sec -280 ft3/sec 280 ft3/sec

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (freckles):

 the base of this box is a L by L in dimensions since it is a square based box now the x there symbolizes the height of the box

OpenStudy (freckles):

The volume of a box=length*height*width

OpenStudy (freckles):

our width=length here

OpenStudy (freckles):

\[V=L^2 x\]

OpenStudy (freckles):

the first sentence tells us what about L'?

OpenStudy (anonymous):

l increases at 2ft/sec

OpenStudy (freckles):

yeah so we are given L'=2ft/sec

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and x'=-2ft/sec

OpenStudy (freckles):

cool

OpenStudy (freckles):

and we are looking for V'

OpenStudy (freckles):

So if V=L^2*x then V'=?

OpenStudy (freckles):

the first rule you want to apply is product rule

OpenStudy (anonymous):

V'=2L(L')(x)+(x')(L^2)

OpenStudy (freckles):

that is beautiful

OpenStudy (freckles):

and we are given L' and x' and we want to find V' when L=12 and x=10

OpenStudy (freckles):

so we just plug in

OpenStudy (freckles):

\[V'=2(12 ft)(2 \frac{ft}{\sec})(10 ft)+(-2 \frac{ft}{\sec})(12 ft)^2\]

OpenStudy (anonymous):

V'=280

OpenStudy (anonymous):

okay thanks I was just a little overwhelmed when I thought there were a lot of variables

OpenStudy (freckles):

\[V'=480 \frac{ft^3}{\sec}-288 \frac{ft^3}{\sec}=(480-288) \frac{ft^3}{\sec}\] is that what you did?

OpenStudy (freckles):

I got a different number for V' but if i made a mistake please correct me

OpenStudy (freckles):

lol

OpenStudy (freckles):

I see I put L=12 and x=10 it was the other way around

OpenStudy (freckles):

\[V'=2(10)(2)(12)-2(10)^2 \\ V'=480-200\]

OpenStudy (freckles):

totally agree with your answer :)

OpenStudy (anonymous):

V'=2L(L')(x)+(x')(L^2) L=10 L'=2 x=12 x'=-2 V'=20(2)(12)+(-2)(100) V'=480-200=280

OpenStudy (anonymous):

ya!

OpenStudy (freckles):

not too terrible right? :) just differentiating and pluggin in :)

OpenStudy (freckles):

anyways have a good day